OpenStudy (anonymous):

Write the following proportions as equivalent fractions. Do not solve If a recipe calls for 1 tsp of vinegar for every cup of milk, how many tsp of vinegar would be required for three cups of milk?

OpenStudy (anonymous):

Okay so 1 tsp of vinegar per 1 cup of milk. It's a proportion: 1:1 1/1 or\[\frac{ 1 }{ 1 }\]

OpenStudy (anonymous):

so would it be, 1:1 x/3?

OpenStudy (anonymous):

\[Solve:\frac{ 1 }{ 1 } \to \frac{ x }{ 3}\]

OpenStudy (anonymous):

okay, i'm not supposed to solve it, but state the equations. Thanks

OpenStudy (anonymous):

welcome(: